the distance between the lines 4x+3y=7 and 8x+6y=15 is

Grade:11

1 Answers

Lakshay Kumar
6 years ago
4x+3y=7.....(1)8x+6y=15...(2)Divide equation (2) by 24x+3y=15/2...(3)From (1) and (3)It`s clear that lines for 4x+3y=7 and 8x+6y=15 are parallelNow applying formula for distance between two parallel lines i.e.,|{(C2-C1)/√(a^2+b^2)}|On applying this|{(15/2-7)/√(16+9)}|=(1/2)/5=5/2 units or 2.5 unitsHope you understand this!!!